### v1

@@ 6.5 SFR: Choice of the models & photometry
In this Section, we validate how the SFRs are affected by the aforementioned choices in the SED framework. The SFR derived via SED fitting (SFR(SED)) depends on the wavelength coverage of the SED fits (e.g. the lack of FIR observations may alter its estimations (Ciesla et al., 2015). Thus, we use an alternative SFR indicator based on the H line luminosity () following the definition given by Kennicutt (1998):
@@
where is corrected for the dust extinction. In this work, we rely on the Balmer decrement method, which is commonly used to correct the line measurements for dust extinction through the comparison of the observed ratio of the H and H emission lines () with the dust-free theoretical value (; Groves et al. 2012). For the theoretical ratio, we used the case B (the optical thick limit) recombination corresponding to a temperature K and electron density (Osterbrock, 1989). The H/H ratio is linked with the attenuation at optical wavelengths by:
@@
For the representative sample (see App. D.1 for the description of the representative sample), we select galaxies with high SNR (SNR ) of H and H line fluxes taken from FastSpecFit v3.2 and correct the H for extinction following:
We also apply aperture correction from FastSpecFit v3.2. Based on the corrected H flux and redshift, we derive the H luminosity and finally, we derive the SFR derived based on the H emission line (SFR(H)) following Eq. 7. To preserve the consistency of the IMF with the SFR definition proposed by Kennicutt (1998), we compare the SFR estimated with CIGALE assuming Salpeter (1955) IMF (see Table 7 for the bias introduced by assumption of the Salpeter (1955) or Chabrier (2003) IMF). We show the comparison of the SFR(SED) and SFR(H) in Fig. 14 for the entire representative sample as well as for star-forming galaxies and AGN. The discrepancy between SFR(SED) and SFR(H) given by and NMAD (see Sec. 5.1 for the definitions of the metrics) are mostly driven by the AGN population. SFRs for star-forming galaxies are in closer agreement ( = 0.136 and NMAD = 0.413) than for AGN ( and NMAD = 0.871). Table 6 shows how the and NMAD changes with redshift. There are no clear trends, the consistency of the SFR(SED) and SFR(H) among different redshift bins suggests that there is no dependence on redshift. The consistency of SFR(SED) and SFR(H) for star-forming galaxies highlights the consistency of the SFRs derived based on the optical-MIR SEDs with the ones derived based on the emission lines.
The degree of the difference in our SFRs estimates from the ones derived by other prescriptions are quantified in Table 7 (see Sec. 5.1 for the definition of the metric). For the star-forming galaxies, the bias introduced by the assumptions of the dust laws or AGN contribution does not affect significantly the SFR estimates ( and NMAD 0.1). The moderate differences ( and NMAD 0.3) are introduced by assumptions over IMF, SSP models, Z, or excluding W3 and W4 from SED fits, while SFH or SED fits obtained solely on optical bands introduce large median difference and NMAD ( and NMAD 0.3). The agreement between SFR(SED) and SFR(H) degrades for AGN, preserving only small and NMAD for the choice of the AGN model ( and NMAD 0.1), and introducing large median difference and NMAD ( and NMAD 0.3) for SFH module and incorporating MIR information to SED fits. The rest of the assumptions introduce a moderate median difference.
@@
| IMF (Chabrier 2003 vs Salpeter 1955) | -0.212 | 0.314 | -0.198 | 0.292 | -0.209 | 0.309 |
| SSP (Bruzual & Charlot 2003 vs Maraston 2005) | -0.214 | 0.462 | -0.256 | 0.384 | -0.227 | 0.389 |
@@
| Z (metallicity fixed to solar vs variable) | 0.217 | 0.598 | 0.064 | 0.146 | 0.163 | 0.374 |
| DustAtt (Calzetti et al. 2000 vs Charlot & Fall 2000) | 0.047 | 0.171 | 0.012 | 0.096 | 0.052 | 0.198 |
| DustEm (Draine et al. 2014 vs Dale et al. 2014) | -0.009 | 0.097 | -0.007 | 0.051 | -0.003 | 0.184 |
| AGN (Fritz et al. 2006 vs no AGN) | -0.142 | 0.220 | -0.049 | 0.075 | -0.190 | 0.287 |
| AGN (Fritz et al. 2006 vs Stalevski et al. 2012, 2016) | 0.008 | 0.038 | 0.002 | 0.010 | 0.044 | 0.127 |
| SEDs ( vs ) | 0.023 | 0.481 | 0.034 | 0.180 | 0.321 | 0.928 |
| SEDs ( vs ) | -0.540 | 1.022 | -0.302 | 0.542 | -0.298 | 0.841 |
| SFR (SED(Salpeter 1955) vs H) | 0.119 | 0.567 | 0.136 | 0.413 | 0.248 | 0.871 |
@@ 6.6 AGN features
As shown in the previous Sections, the incorporation of the WISE photometry has an influence on the stellar mass and SFR estimates (see Sec. 6.4 and 6.5). Here, we explore how the AGN features derived with CIGALE depend on the inclusion of the MIR photometry in SED fits. The change of the AGN fraction (AGNFRAC) and viewing angle (AGNPSY) for the BL AGN, NL AGN and star-forming galaxies drawn from the representative sample (see App. D.1 for a description of the representative sample) when using solely optical information (grz) or optical and WISE W1 and W2 (grzW12) or optical and all WISE (grzW14) is shown in Fig. 15 and 16, respectively. As expected, optical SED fits cannot distinguish the contribution of AGN to IR, which results in an artificial AGN fraction (AGNFRAC ) and viewing angle (AGNPSY ) for almost all galaxies independent of their galaxy type. The AGNFRAC and AGNPSY distributions start to differentiate for BL AGN, NL AGN, and star-forming galaxies when at least W12 are included in SED fitting. When the W12 is incorporated in the SED fits, the BL AGN are characterized by lower AGNFRAC than NL AGN (with a median AGNFRAC increasing from 0.16 to 0.24) while star-forming galaxies have an AGNFRAC distribution similar to the one for NL AGN (with a median 0.22). Interestingly, when including W12, the AGNPSY starts to properly identify BL AGN based on their high viewing angle (AGNPSY ), although there is also a strong peak at AGNPSY . Both star-forming and NL AGN are characterized by similar viewing angles (AGNPSY ).
@@
The incorporation of all four WISE bands () results in smoothing the AGNFRAC distribution for BL AGN with a strong peak at AGNFRAC and AGNPSY , while for NL AGN and star-forming galaxies the distribution is characterized by a strong peak at AGNFRAC and a secondary peak at . The difference in AGNFRAC between star-forming and NL AGN is in the long tail towards high values for NL AGN, while the fraction of star-forming galaxies exceeding AGNFRAC is negligible. The AGNPSY distributions for NL AGN and star-forming galaxies are similarly characterized by wide peaks at AGNPSY , however, a small fraction is characterized by the viewing angle typical for BL AGN. Even with MIR SED fits the distribution of AGNFRAC and AGNPSY for NL AGN and star-forming galaxies are similar, challenging the usefulness of these AGN features to differentiate between AGN and non-AGN. Nevertheless, BL AGN show clear preference for AGNFRAC and AGNPSY .
@@
Independently on the availability of the WISE photometry, 80% of BL AGN are identified as AGN based on the AGNFRAC . Following the same criterion, 65% of NL AGN are also classified as AGN. A similar fraction (68%) of star-forming galaxies is characterized by AGNFRAC , but is reduced to 19% if all WISE bands with SNR are included in the SED fits (FLAGINFRARED = 4). This confirms that without the MIR information, the star-forming galaxies overestimate the AGN contribution, although providing measurements in all four WISE bands reduces significantly the fraction of the star-forming galaxies with overestimated AGN fraction (from 68% to 19%).
@@
The physical properties derived for the majority (68%) of star-forming galaxies are thus derived under the assumption of the false presence of AGN for sources without the MIR information. However, we validate that the stellar mass estimates can still be securely used. For a sample of star-forming galaxies with AGNFRAC , we compare the derived stellar masses with the ones estimated without incorporating AGN contribution (i.e. the AGNFRAC is fixed to the null value) independently of the availability of the WISE photometry. Both the and NMAD are comparable (with = 0.014 and NMAD = 0.037). Similarly, the effect on the SFR is negligible ( = -0.079 and NMAD = 0.119). For star-forming galaxies with a low AGN fraction (AGNFRAC ), there is no impact on stellar mass or SFR estimates (with and NMAD = 0.011, 0.025 for stellar mass and SFR, respectively). To summarize, although the AGN features especially for sources lacking high SNR WISE photometry are not reliable, the overestimation of the AGN fraction for star-forming galaxies does not affect their stellar mass or SFR estimates.
@@ 7 Summary
We created the VAC of physical properties (including stellar masses, SFRs, rest-framed magnitudes, and AGN features) for DESI EDR galaxies. Here, we briefly summarize the VAC:
@@
The catalog includes almost 1.3 million galaxies with optical and NIR photometry (grz) complemented by WISE (W14) information spanning over a wide redshift range (; for details about data see Sec. 2).
@@
Several quality flags, including the quality of the fit expressed by or the PDF of the derived properties, or the SNR of available photometry, among others, allow the users to adapt the sample to their scientific goals (see Sec. 4.1).
@@
We find very good agreement with COSMOS and other VACs that are BC03-based. The relatively good agreement with the SDSS(Firefly) VAC which is obtained by fitting spectra and not photometry and using another model prescription (SSP, SFH, metallicities) showcases the excellent agreement of the stellar mass estimates with other VACs (see Sec. 5 and Tab. 4).
@@
The presented VAC highlights the importance of incorporating MIR photometry in the SED fits. Without WISE photometry the AGN fraction is overestimated for star-forming galaxies, though it does not affect their stellar mass or SFR estimates (see Sec. 6.6). We show that accounting for AGN contribution is essential in deriving stellar masses for galaxies hosting AGN. With CIGALE we are able to recover both stellar masses and AGN luminosities consistent with other methods. The host properties of the DESI AGN will be explored in detail in future work.
Acknowledgements
The authors thank the anonymous referee for insightful comments. The authors would like to express their appreciation to Laure Ciesla for sharing the CIGALE non-parametric SFH module sfhNlevels, J. Thorne, and H. Suh for sharing their catalogs.
This work has been supported by the Polish National Agency for Academic Exchange (Bekker grant BPN/BEK/2021/1/00298/DEC/1) and the European Union’s Horizon 2020 Research and Innovation program under the Maria Sklodowska-Curie grant agreement (No. 754510). M.S. acknowledges financial support from the State Research Agency of the Spanish Ministry of Science and Innovation under the grants ’Galaxy Evolution with Artificial Intelligence’ with reference PGC2018-100852-A-I00 and ’BASALT’ with reference PID2021-126838NB-I00. M.M. acknowledges support from the Spanish Ministry of Science and Innovation through the project PID2021-124243NB-C22. This work was partially supported by the program Unidad de Excelencia María de Maeztu CEX2020-001058-M. H.Z. acknowledges the support from the National Natural Science Foundation of China (NSFC; grant Nos. 12120101003 and 12373010) and National Key R&D Program of China (grant Nos. 2023YFA1607800, 2022YFA1602902) and Strategic Priority Research Program of the Chinese Academy of Science (Grant Nos. XDB0550100).
This material is based upon work supported by the U.S. Department of Energy (DOE), Office of Science, Office of High-Energy Physics, under Contract No. DE–AC02–05CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences under Contract No. AST-0950945 to the NSF’s National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Humanities, Science and Technology of Mexico (CONAHCYT); the Ministry of Science, Innovation and Universities of Spain (MICIU/AEI/10.13039/501100011033), and by the DESI Member Institutions: https://www.desi.lbl.gov/collaborating-institutions. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U. S. National Science Foundation, the U. S. Department of Energy, or any of the listed funding agencies.
The authors are honored to be permitted to conduct scientific research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation.
The DESI Legacy Imaging Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS), the Beijing-Arizona Sky Survey (BASS), and the Mayall z-band Legacy Survey (MzLS). DECaLS, BASS and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSF’s NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. NOIRLab is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. Pipeline processing and analyses of the data were supported by NOIRLab and the Lawrence Berkeley National Laboratory. Legacy Surveys also uses data products from the Near-Earth Object Wide-field Infrared Survey Explorer (NEOWISE), a project of the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration. Legacy Surveys was supported by: the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy; the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility; the U.S. National Science Foundation, Division of Astronomical Sciences; the National Astronomical Observatories of China, the Chinese Academy of Sciences and the Chinese National Natural Science Foundation. LBNL is managed by the Regents of the University of California under contract to the U.S. Department of Energy. The complete acknowledgments can be found at https://www.legacysurvey.org/.
Data Availability
The VAC of physical properties of DESI galaxies is publicly available at https://data.desi.lbl.gov/doc/releases/edr/vac/cigale/. The presented analysis is conducted on v1.2, while the latest available version v1.4 includes an additional 51,126 () that were missing information about the RELEASE in the previous version. There is also a supplementary table of 8422 sources for which physical properties are based on the redshift coming from the QSO afterbruner pipeline relying on the QuasrNet and the broad Mg II finder pipelines. The afterburner shows better performance (94%) in identifying visually inspected QSOs relative to Redrock pipeline (86%; see Chaussidon et al. 2023 and Alexander et al. 2023 for more details). The VAC for the forthcoming release (DESI Year 1) will be published as part of Data Release 1 similar to the currently public DESI EDR.

### v2

@@ 6.5 SFR: Choice of the models and photometry
In this section, we validate how the SFRs are affected by the aforementioned choices in the SED framework. The SFR derived via SED fitting (SFR(SED)) depends on the wavelength coverage of the SED fits (e.g., the lack of FIR observations may alter its estimations (Ciesla et al., 2015). Thus, we used an alternative SFR indicator based on the H line luminosity (), following the definition given by Kennicutt (1998):
@@
where is corrected for the dust extinction. In this work, we rely on the Balmer decrement method, which is commonly used to correct the line measurements for dust extinction through the comparison of the observed ratio of the H and H emission lines () with the dust-free theoretical value (; Groves et al. 2012). For the theoretical ratio, we used the case B (the optical thick limit) recombination, corresponding to a temperature of K and electron density of (Osterbrock, 1989). The H/H ratio is linked with the attenuation at optical wavelengths by:
@@
For the representative sample (see Appendix D.1 for the description of the representative sample), we selected galaxies with high S/N (S/N ) of H and H line fluxes taken from FastSpecFit v3.2 and correct the H for extinction following:
We also applied an aperture correction from FastSpecFit v3.2. Based on the corrected H flux and redshift, we derived the H luminosity; finally, we derived the SFR derived based on the H emission line (SFR(H)) following Eq. 7. To preserve the consistency of the IMF with the SFR definition proposed by Kennicutt (1998), we compared the SFR estimated with CIGALE assuming Salpeter (1955) IMF . We refer to Table 7 for the bias introduced by assumption of the Salpeter (1955) or Chabrier (2003) IMF). We show the comparison of the SFR(SED) and SFR(H) in Fig. 14 for the entire representative sample as well as for star-forming galaxies and AGNs. The discrepancy between SFR(SED) and SFR(H) given by and NMAD (see Sect. 5.1 for the definitions of the metrics) are mostly driven by the AGN population. The SFRs for star-forming galaxies are in closer agreement ( = 0.136 and NMAD = 0.413) than for AGNs ( and NMAD = 0.871). Table 6 shows how the and NMAD changes with redshift. There are no clear trends, the consistency of the SFR(SED) and SFR(H) among different redshift bins suggests that there is no dependence on redshift. The consistency of SFR(SED) and SFR(H) for star-forming galaxies highlights the consistency of the SFRs derived based on the optical-MIR SEDs with the ones derived based on the emission lines.
The degree of the difference in our SFRs estimates from the ones derived by other prescriptions are quantified in Table 7 (see Sect. 5.1 for the definition of the metric). For the star-forming galaxies, the bias introduced by the assumptions of the dust laws or AGN contribution does not significantly affect the SFR estimates ( and NMAD 0.1). The moderate differences ( and NMAD 0.3) are introduced by assumptions over IMF, SSP models, Z, or excluding W3 and W4 from SED fits, while the SFH or SED fits obtained solely on optical bands introduce large median differences and NMAD ( and NMAD 0.3). The agreement between SFR(SED) and SFR(H) degrades for AGNs, preserving only small and NMAD for the choice of the AGN model ( and NMAD 0.1), while introducing large median difference and NMAD ( and NMAD 0.3) for SFH module and incorporating MIR information to SED fits. The rest of the assumptions introduced a moderate median difference.
@@
| IMF (Chabrier 2003 vs. Salpeter 1955) | -0.212 | 0.314 | -0.198 | 0.292 | -0.209 | 0.309 |
| SSP (Bruzual & Charlot 2003 vs. Maraston 2005) | -0.214 | 0.462 | -0.256 | 0.384 | -0.227 | 0.389 |
@@
| Z (metallicity fixed to solar vs. variable) | 0.217 | 0.598 | 0.064 | 0.146 | 0.163 | 0.374 |
| DustAtt (Calzetti et al. 2000 vs. Charlot & Fall 2000) | 0.047 | 0.171 | 0.012 | 0.096 | 0.052 | 0.198 |
| DustEm (Draine et al. 2014 vs. Dale et al. 2014) | -0.009 | 0.097 | -0.007 | 0.051 | -0.003 | 0.184 |
| AGN (Fritz et al. 2006 vs. no AGN) | -0.142 | 0.220 | -0.049 | 0.075 | -0.190 | 0.287 |
| AGN (Fritz et al. 2006 vs. Stalevski et al. 2012, 2016) | 0.008 | 0.038 | 0.002 | 0.010 | 0.044 | 0.127 |
| SEDs ( vs. ) | 0.023 | 0.481 | 0.034 | 0.180 | 0.321 | 0.928 |
| SEDs ( vs. ) | -0.540 | 1.022 | -0.302 | 0.542 | -0.298 | 0.841 |
| SFR (SED(Salpeter 1955) vs. H) | 0.119 | 0.567 | 0.136 | 0.413 | 0.248 | 0.871 |
@@ 6.6 AGN features
As shown in the previous sections, the incorporation of the WISE photometry has an influence on the stellar mass and SFR estimates (see Sect. 6.4 and 6.5). Here, we explore how the AGN features derived with CIGALE depend on the inclusion of the MIR photometry in SED fits. The change of the AGN fraction (AGNFRAC) and viewing angle (AGNPSY) for the BL AGNs, NL AGNs, and star-forming galaxies drawn from the representative sample (see Appendix D.1 for a description of the representative sample) when using 1) solely optical information (grz), 2) optical and WISE W1 and W2 (grzW12), or 3) optical and all WISE (grzW14) is shown in Figs. 15 and 16, respectively. As expected, optical SED fits cannot distinguish the contribution of AGN to IR, which results in an artificial AGN fraction (AGNFRAC ) and viewing angle (AGNPSY ) for almost all galaxies independent of their galaxy type. The AGNFRAC and AGNPSY distributions start to differentiate for BL AGNs, NL AGNs, and star-forming galaxies when at least W12 is included in SED fitting. When the W12 is incorporated in the SED fits, the BL AGNs are characterized by lower AGNFRAC than NL AGNs (with a median AGNFRAC increasing from 0.16 to 0.24); whereas star-forming galaxies have an AGNFRAC distribution similar to that of NL AGN (with a median 0.22). Interestingly, when including W12, AGNPSY starts to properly identify BL AGNs based on their high viewing angle (AGNPSY ), although there is also a strong peak at AGNPSY . Both star-forming and NL AGNs are characterized by similar viewing angles (AGNPSY ).
@@
The incorporation of all four WISE bands () results in smoothing the AGNFRAC distribution for BL AGNs with a strong peak at AGNFRAC and AGNPSY ; whereas for NL AGNs and star-forming galaxies, the distribution is characterized by a strong peak at AGNFRAC and a secondary peak at . The difference in AGNFRAC between star-forming and NL AGNs is in the long tail towards high values for NL AGNs, while the fraction of star-forming galaxies exceeding AGNFRAC is negligible. The AGNPSY distributions for NL AGNs and star-forming galaxies are similarly characterized by wide peaks at AGNPSY ; however, a small fraction is characterized by the viewing angle typical for BL AGNs. Even with MIR SED fits, the distribution of AGNFRAC and AGNPSY for NL AGNs and star-forming galaxies are similar, challenging the usefulness of these AGN features to differentiate between AGN and non-AGN galaxies. Nevertheless, BL AGNs show a clear preference for AGNFRAC and AGNPSY .
@@
Independently of the availability of WISE photometry, 80% of BL AGNs are identified as AGNs based on the AGNFRAC . Following the same criterion, 65% of NL AGNs are also classified as AGNs. A similar fraction (68%) of star-forming galaxies is characterized by AGNFRAC , but this is reduced to 19% if all WISE bands with S/N are included in the SED fits (FLAGINFRARED = 4). This confirms that without the MIR information, the star-forming galaxies overestimate the AGN contribution, although providing measurements in all four WISE bands reduces significantly the fraction of the star-forming galaxies with overestimated AGN fraction (from 68% to 19%).
@@
The physical properties derived for the majority (68%) of star-forming galaxies are thus derived under the assumption of the false presence of AGN for sources without the MIR information. However, we validated the notion that the stellar mass estimates can still be securely used. For a sample of star-forming galaxies with AGNFRAC , we compared the derived stellar masses with the ones estimated without incorporating AGN contribution (i.e., the AGNFRAC is fixed to the null value) independently of the availability of WISE photometry. Both the and NMAD are comparable (with = 0.014 and NMAD = 0.037). Similarly, the effect on the SFR is negligible ( = -0.079 and NMAD = 0.119). For star-forming galaxies with a low AGN fraction (AGNFRAC ), there is no impact on stellar mass or SFR estimates (with and NMAD = 0.011, 0.025 for stellar mass and SFR, respectively). To summarize, although the AGN features especially for sources lacking high S/N WISE photometry are not reliable, the overestimation of the AGN fraction for star-forming galaxies does not affect their stellar mass or SFR estimates.
@@ 7 Summary
We created the VAC of physical properties (including stellar masses, SFRs, rest-frame magnitudes, and AGN features) for DESI EDR galaxies. Here, we briefly summarize the VAC:
@@
The catalog includes almost 1.3 million galaxies with optical and NIR photometry (grz) supplemented with WISE (W14) information spanning a broad redshift range (; for details, see Sect. 2).
The physical properties are derived with the CIGALE SED fitting code (Boquien et al. 2019) including AGN templates from Fritz et al. 2006 (see Sect. 3 and Table 3 for details on the SED fitting prescription). Tables 5 and 7 elaborate on the extent to which our estimates depend on our model assumptions.
@@
Several quality flags, such as the quality of the fit expressed by , the PDF of the derived properties, or the S/N of available photometry, would allow users to adapt the sample to specific scientific goals (see Sect. 4.1).
@@
We found a very good agreement with COSMOS and other VACs that are BC03-based. The relatively good agreement with the SDSS(Firefly) VAC, which is obtained by fitting the spectra, rather than the photometry. Using another model prescription (SSP, SFH, metallicities, etc.) showcases the excellent agreement of the stellar mass estimates with other VACs (see Sect. 5 and Table 4).
@@
The VAC presented here highlights the importance of incorporating MIR photometry in the SED fits. Without the inclusion of WISE photometry, the AGN fraction would be overestimated for star-forming galaxies, although it does not affect their stellar mass or SFR estimates (see Sect. 6.6). We demonstrate that accounting for the AGN contribution is essential for deriving stellar masses for galaxies hosting AGN. With CIGALE, we are able to recover both stellar masses and AGN luminosities in a way that is consistent with other methods. We note that the particular host properties of DESI AGNs will be explored in detail in a future work.
Acknowledgements.
The presented analysis is conducted on the VAC of physical properties of DESI galaxies v1.2, while the latest available version v1.4 includes an additional 51,126 () that were missing information about the RELEASE in the previous version. There is also a supplementary table of 8,422 sources for which physical properties are based on the redshift coming from the QSO afterbruner pipeline relying on the QuasrNet and the broad Mg II finder pipelines. The afterburner shows better performance (94%) in identifying visually inspected QSOs relative to Redrock pipeline (86%; see Chaussidon et al. 2023 and Alexander et al. 2023 for more details). The VAC for the forthcoming release (DESI Year 1) will be published as part of Data Release 1 similar to the currently public DESI EDR. The authors thank the anonymous referee for insightful comments. The authors would like to express their appreciation to Laure Ciesla for sharing the CIGALE non-parametric SFH module sfhNlevels, J. Thorne, and H. Suh for sharing their catalogs. This work has been supported by the Polish National Agency for Academic Exchange (Bekker grant BPN/BEK/2021/1/00298/DEC/1) and the European Union’s Horizon 2020 Research and Innovation program under the Maria Sklodowska-Curie grant agreement (No. 754510). M.S. acknowledges financial support from the State Research Agency of the Spanish Ministry of Science and Innovation under the grants ’Galaxy Evolution with Artificial Intelligence’ with reference PGC2018-100852-A-I00 and ’BASALT’ with reference PID2021-126838NB-I00. M.M. acknowledges support from the Spanish Ministry of Science and Innovation through the project PID2021-124243NB-C22. This work was partially supported by the program Unidad de Excelencia María de Maeztu CEX2020-001058-M. H.Z. acknowledges the support from the National Natural Science Foundation of China (NSFC; grant Nos. 12120101003 and 12373010) and National Key R&D Program of China (grant Nos. 2023YFA1607800, 2022YFA1602902) and Strategic Priority Research Program of the Chinese Academy of Science (Grant Nos. XDB0550100). This material is based upon work supported by the U.S. Department of Energy (DOE), Office of Science, Office of High-Energy Physics, under Contract No. DE–AC02–05CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences under Contract No. AST-0950945 to the NSF’s National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Humanities, Science and Technology of Mexico (CONAHCYT); the Ministry of Science, Innovation and Universities of Spain (MICIU/AEI/10.13039/501100011033), and by the DESI Member Institutions: https://www.desi.lbl.gov/collaborating-institutions. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U. S. National Science Foundation, the U. S. Department of Energy, or any of the listed funding agencies. The authors are honored to be permitted to conduct scientific research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation. The DESI Legacy Imaging Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS), the Beijing-Arizona Sky Survey (BASS), and the Mayall z-band Legacy Survey (MzLS). DECaLS, BASS and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSF’s NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. NOIRLab is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. Pipeline processing and analyses of the data were supported by NOIRLab and the Lawrence Berkeley National Laboratory. Legacy Surveys also uses data products from the Near-Earth Object Wide-field Infrared Survey Explorer (NEOWISE), a project of the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration. Legacy Surveys was supported by: the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy; the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility; the U.S. National Science Foundation, Division of Astronomical Sciences; the National Astronomical Observatories of China, the Chinese Academy of Sciences and the Chinese National Natural Science Foundation. LBNL is managed by the Regents of the University of California under contract to the U.S. Department of Energy. The complete acknowledgments can be found at https://www.legacysurvey.org/.Data availability